0.0.2702 QUESTION 17 Consider the following partial ANOVA table. Source of variation df Sum of squares Mean squares Treatments Error Total 6.67 75 60 19 135 25 3.75 The numerator and denominator degrees of freedom (identified by asterisks) are, respective 1. 4 and 15 2. 3 and 16 3. 15 and 4 4. 16 and 3 5. 4 and 8
Answer #1

df =SS/MS

for numerator

df = 75/25 = 3

for denominator

df = 60/3.75 = 16

hence

option 2) 3 and 16 is correct

Please rate